an ideal gas of monotonic molecules expands from its initial state A to a state B through an isobaric process and then further expands to a volume C . Find the work done by the gas ,increase in internal energy, and the energy transferred by heat to the gas over the whole process


1
Expert's answer
2021-06-14T17:27:09-0400

"A_{AB}=p_1\Delta V=4\cdot 1.49=5.96~J,"

"A_{BC}=\frac{p_2V_2}{\gamma -1}(1-(\frac{V_2}{V_3})^{\gamma -1})=\frac{4\cdot 3.49}{2/3}\cdot(1-(\frac{3.49}{8})^{2/3})=8.90~J,"

"A=5.96+8.90=14.86~J,"


"\Delta U_{AB}=\frac 32 p_1 \Delta V=8.94~J,"

"\Delta U_{BC}=-A_{BC}=-8.90~J,"

"\Delta U=8.94-8.90=0.04~J,"


"Q_{AB}=A_{AB}+\Delta U_{AB}=5.96+8.94=14.90~J,"

"Q_{BC}=0~J,"

"Q=14.90+0=14.90~J."